jalissa50925:

A cone has a height of 16 centimeters and a radius of 12 centimeters. What is the exact lateral and surface area of the cone? Type the correct answer in each box. Use numerals instead of words. The lateral area is square centimeters. The total surface area is square centimeters.

jalissa50925:

thank you

surjithayer:

r=radius of cone l=slant height h=height \[base~area=\pi r^2=\pi (12)^2=144 \] \[l=\sqrt{r^2+h^2}=\sqrt{12^2+16^2}=\sqrt{144+256}=\sqrt{400}=20~cm\]

surjithayer:

r=radius of cone l=slant height h=height \[base~area=\pi r^2=\pi (12)^2=144 \pi \] \[l=\sqrt{r^2+h^2}=\sqrt{12^2+16^2}=\sqrt{144+256}=\sqrt{400}=20~cm\]

surjithayer:

\[slant~area=\pi r l=\pi \times 12\times 20=240 \pi\] \[total~ area=144 ~\pi+240 ~\pi=384~\pi\] \[\approx1206.37 cm^2\]